# Principled Federated Random Forests for Heterogeneous Data

### v1

@@ D.2 Covariate and Outcome Shift
Figure 6 shows that our federated FedForest Quantile and AvgImp methods that split on $H$ perform comparatively to their centralized counterpart.
D.3 Heterogeneity diagnostics for adaptive optimization
@@
We test for covariate shift by assessing the null hypothesis $H_{0}:X\perp\!\!\!\perp H$ using a federated discriminator of clients. Instead of estimating density ratios—either parametrically or via kernel methods with high communication cost—we reuse the impurity-based split selection of federated decision trees by treating the client index $H$ as the target variable. A federated Random Forest is trained to predict $H$ from $X$ using standard CART criteria.
@@
$\Delta{I}_{H}(j,t;\mathcal{S}_{\nu})\!\!=\!I_{\text{site}}(\mathcal{S}_{\nu})-\left(\frac{n_{L}}{n_{\nu}}I_{H}(\mathcal{S}_{L})+\frac{n_{R}}{n_{\nu}}I_{H}(\mathcal{S}_{R})\right),$

### v2

@@ F.2 Covariate and outcome shifts
The results align with previous conclusions from the covariate shift and outcome shift paragraphs in Section 5.
F.3 Robustness to outliers
To assess robustness to outliers, we use the same data-generating process as in the homogeneous (i.i.d.) simulation setting described in Section 5. We then contaminate the training data by introducing outliers at fractions ranging from $0\%$ to $20\%$ of the total sample size. For the contaminated observations, the corresponding feature vectors are multiplied by a factor of $1000$, producing extreme covariate values while leaving the remaining data unchanged.
Table 7: Model Degradation (MSE) vs. Outlier Fraction. Results show the mean $\pm$ standard deviation across 2020 simulations.
Outlier Fraction
Method 0% 2% 5% 10% 15% 20%
Centralized Scikit X $0.48\pm 0.05$ $0.48\pm 0.05$ $0.48\pm 0.07$ $0.52\pm 0.10$ $0.56\pm 0.09$ $0.53\pm 0.08$
Fed. Average Impurity X (ours) $1\pm 0.12 0.55±0$$5\pm 0.09$ $0.63\pm 0.08$ $0.64\pm 0.11$ $5\pm 0.11 0.66±0$$6\pm 0.09$
Fed. Quantiles X (ours) $0.46\pm 0.08$ $0.48\pm 0.07$ $0.55\pm 0.08$ $0.54\pm 0.09$ $0.59\pm 0.09$ $0.56\pm 0.08$
Fed. Histogram (Kalloori et al.) $0.51\pm 0.11$ $0.53\pm 0.08$ $0.67\pm 0.15$ $0.90\pm 0.18$ $1.09\pm 0.32$ $1.55\pm 0.21$
Fed. Ensemble (Hauschild et al.) $1.10\pm 0.37$ $1.08\pm 0.23$ $1.33\pm 0.24$ $1.30\pm 0.32$ $1.47\pm 0.31$ $1.41\pm 0.17$
Local Learning $1.27\pm 0.41$ $1.22\pm 0.24$ $1.45\pm 0.27$ $1.44\pm 0.34$ $1.61\pm 0.35$ $1.55\pm 0.21$
The results in Table 7 show that the proposed methods are substantially more stable under outlier contamination than the histogram-based federated baseline. In particular, the MSE of Fed. Histogram increases sharply as the outlier fraction grows, rising from $0.51$ with no contamination to $1.55$ at $20\%$ contamination. By contrast, Fed. Quantiles remains close to the centralized reference across all contamination levels, with MSE increasing only from $0.46$ to $0.56$. Fed. Average Impurity also degrades gradually, reaching $0.66$ at the highest contamination level.
Histogram-based methods are sensitive to outliers because global min–max binning can stretch the feature range, collapsing most clean samples into coarse bins and reducing split quality. In contrast, quantile-based methods rely on ranks, so extreme values are confined to tail bins while the remaining bins preserve resolution over the clean data. This accounts for their milder degradation under outlier contamination.
F.4 Heterogeneity diagnostics for adaptive optimization
@@
We test for covariate shift by assessing the null hypothesis $H_{0}:X\perp\!\!\!\perp H$ using a federated discriminator of clients. Instead of estimating density ratios—either parametrically or via kernel methods with high communication cost—we reuse the impurity-based split selection of federated decision trees by treating the client index $H$ as the target variable. A federated Random Forest is trained to predict $H|X$ using standard CART criteria.
@@
$\Delta{I}_{H}(j,t;\mathcal{S}_{\nu})\!\!=\!I_{H}(\mathcal{S}_{\nu})-\left(\frac{n_{L}}{n_{\nu}}I_{H}(\mathcal{S}_{L})+\frac{n_{R}}{n_{\nu}}I_{H}(\mathcal{S}_{R})\right),$